DOPPLER-FREE TWO-PHOTON ABSORPTION EXCITATION SPECTROSCOPY AND THE ZEEMAN EFFECT OF THE A ^{1}B_{2u} ← X ^{1}A_{1g} 14_{0}^{1} BAND OF C_{6}D_{6}

J. WANG, A. DOI, S. KASAHARA, and \underline{HAJIME KATO}, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; M. BABA, Department of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8501, Japan.

A Doppler-free excitation spectrum and the Zeeman effect of the A ^{1}B_{2u}(v_{14} = 1) ← X ^{1}A_{1g}(v = 0) transition of C_{6}D_{6} have been measured by means of two-photon absorption spectroscopy with counter-propagating light beams of identical wavelength within an external cavity. Rotational lines were fully resolved, and 1357 Q_{Q} lines of J = 0−64, K = 0−64 have been assigned, and the molecular constants of the A ^{1}B_{2u}(v_{14} = 1) state have been determined. Perturbations centerd at K = 28 − 29 were observed for J ≥ 30, and these are identified as originating from the perpendicular Colioris interaction. The Zeeman splittings for lines of a given J were observed to increase proportionally to K^{2} and reach a maximum at K = J. This demonstrates that the magnetic moment lies along the c axis (perpendicular to the molecular plane). The magnetic moment of the A ^{1}B_{2u}(v_{14} = 1, J = 64, K = 64) level was determined to be 0.011 μ_{B}. The Zeeman splittings of the K = J levels were observed to increase linearly with J. From the analysis, the magnetic moment is shown to be originating mostly from mixing of the S_{1} ^{1}B_{2u} and S_{2} ^{1}B_{1u} states by the J-L coupling (electronic Coriolis interaction). The Zeeman splitting which could be identified as originating from a singlet-triplet interaction was not observed.